# M3LEO: A Multi-Modal, Multi-Label Earth Observation Dataset Integrating Interferometric SAR and RGB Data

### v1

@@ Appendix B Training Hyperparameters
Hyperparameter details for all experiments can be seen in Table B.1. We did not perform significant hyperparameter tuning. Model selection was using best performance on the validation set. All models used approximately $7.9\times 10^{6}$ trainable parameters, to the nearest $10^{5}$. All experiments were performed on virtual machines provisioned via Google Compute Engine, using 2 NVIDIA V100 GPUs. We show runtimes for each experiment in Table B.2. The required channels from each dataset were cached using Blosc2 to reduce runtime. We do not provide data in this format as it is less able to handle relevant metadata than the .tif and .nc files we provide. We provide tools for performing this caching within our framework.
Table B.1: Hyperparameter details for baseline experiments
@@
Table B.2: Runtimes for single-input and data fusion baseline experiments
@@
S2+GSSIC 19h 23m 68h 20m** 19h 49m
    Intermittent slowdown occured occasionally due to suspected IO issues occurring with specific hardware setups. We aim to confirm and correct the source of these issues in an update to the framework. We emphasise that these issues are occasional.
    Slowdown resulting from reduced number of workers to avoid memory problems resulting from *.
Appendix C ESA World Cover Land Cover Classes
We outline the 11 classes that together comprise the ESAWC dataset in Table C.1, along with their pixel values in M3LEO. We opted to leave these pixel values as-found in the original ESAWC.
Table C.1: ESAWC Classes along with pixel values for data as-found in M3LEO.
@@ D.1 Data Limitations
While the M3LEO dataset is large, coverage is not global. We limited coverage to the area covered in the GUNW dataset, as one aim of this work is to encourage the uptake of interferometric SAR data. The coverage of GUNW, however, is limited to the areas that were of interest to the Alaska Satellite Facility who originally distributed this data. Generating further inteferometric data is substantially technically complex, computationally demanding and requires the use of SAR acquisitions with phase information (which are difficult to access compared to the amplitude data we provide). Were this data to be generated, cloud storage requirements for M3LEO would approach the petabyte scale, for which we are unable to provide a feasible long-term storage solution.
@@
We highlight the potential application of interferometric data to change detection tasks, but note that this data is not included in the initial relase of M3LEO. Two datasets that could be used for change detection tasks have been processed - namely, ESA CCI Burned Area MODIS [^11] https://developers.google.com/earth-engine/datasets/catalog/ESA_CCI_FireCCI_5_1 and the Global Flood Database [^12] https://developers.google.com/earth-engine/datasets/catalog/GLOBAL_FLOOD_DB_MODIS_EVENTS_V1 - but have not been tested extensively enough for inclusion here. We are unable to guarantee the release of these components simultaneously with the data advertised in the main text of this work, but aim to release them in the future.
@@
Data for all input datasets is provided for 2020 only in the initial release. Data from S1GRD, S2RGB has been tiled for 2018-2021 additionally, but is not provided in the initial release due to the large storage requirements for input datasets, and the lack of processed interferometric/coherence data for the same years. Despite this, we opted to include multitemporal data for MODISVEG for 2016-2020 as the storage requirements were far smaller.
D.2 Framework Limitations
@@
Data is currently loaded from the disk using xarray. We chose to use xarray as it easily accommodates handling the wealth of metadata associated with remote sensing imagery, but it is not performant for loading a large number of tiles quickly. For users who wish to access the same data many times - either over many epochs, or many experiments - we recommend caching the returned data arrays at the first encounter. We provide this facility using blosc2. The caching process can be computationally expensive for large datasets, but is relatively far cheaper than performing repeat runs using xarray. Slightly increased disk space requirements are incurred.
@@
We provide data chips at a fixed spatial size of $448\times 448$m. While we provide the facility to re-tile data straightforwardly at different sizes with geetiles and sartiles, this process is highly computationally expensive when tiling over large spatial areas.
@@
While we provided some baseline results using our framework and data, we did not provide a benchmarking framework under which models could be compared. For example, we made no assertion as to whether models should be evaluated on chips with missing channels - we chose to fill any missing data with a dummy value of 00. We also did not include European data in our test set. Although comparisons could be made by copying our approach exactly, we encourage domain experts to evaluate models according to the needs of their particular application.
Appendix E Licenses
The licenses under which each component of M3LEO was originally distributed are listed in Table E. We distribute our framework and data under the Creative Commons BY-SA 4.0 license..

### v2

@@ D.3 Training Hyperparameters
Hyperparameter details for supervised experiments can be seen in Table D.2. We did not perform significant hyperparameter tuning. Model selection was using best performance on the validation set. All models used approximately $7.9\times 10^{6}$ trainable parameters, to the nearest $10^{5}$. We show runtimes for each experiment on 2 NVIDIA V100 GPUs in Table D.3.
Table D.2: Hyperparameter details for supervised experiments
@@
Table D.3: Runtimes for single-input and data fusion baseline experiments
@@
S2+GSSIC 19h 23m 68h 20m 19h 49m
Appendix E ESA World Cover Land Cover Classes
We outline the 11 classes that together comprise the ESAWC dataset in Table E.1, along with their pixel values in M3LEO. We opted to leave these pixel values as-found in the original ESAWC.
Table E.1: ESAWC Classes along with pixel values for data as-found in M3LEO.
@@ F.1 Data Limitations
While the M3LEO dataset is large, coverage is not global. We limited coverage to the area covered in the GUNW dataset, which is approximately equal to the regions in which Senintel-1 has dual-polarization, ascending-descending coverage. Generating further inteferometric data is substantially technically complex, computationally demanding and requires the use of SAR acquisitions with phase information (which are difficult to access compared to the amplitude data we provide). Were this data to be generated, cloud storage requirements for M3LEO would approach the petabyte scale, for which we are unable to provide a feasible long-term storage solution.
@@
We highlight the potential application of interferometric data to change detection tasks, but note that this data is not included in the initial relase of M3LEO. Two datasets that could be used for change detection tasks have been processed — namely, ESA CCI Burned Area MODIS121212developers.google.com/earth-engine/datasets/catalog/ESA_CCI_FireCCI_5_1 and the Global Flood Database131313developers.google.com/earth-engine/datasets/catalog/GLOBAL_FLOOD_DB_MODIS_EVENTS_V1 — but have not been tested extensively enough for inclusion here. We are unable to guarantee the release of these components simultaneously with the data advertised in the main text of this work, but aim to release them in the future.
@@
Data from S1GRD, S2SRM and AGB have been tiled for 2018-2020 additionally, but other datasets are provided for 2020 only. Many datasets, such as ESAWC, simply do not exist for 2018 or 2019. One satellite of the Sentinel-1 pair, Sentinel-1B, suffered a power unit failure in December 2021, so we cannot provide data with the same coverage as 2018-2020 from 2021 onwards.
F.2 Framework Limitations
@@
Data is currently loaded from the disk using xarray. We chose to use xarray as it easily accommodates handling the wealth of metadata associated with remote sensing imagery, but it is not performant for loading a large number of tiles quickly. For users who wish to access the same data many times — either over many epochs, or many experiments — we recommend caching the returned data arrays at the first encounter. We provide this facility using blosc2. The caching process can be computationally expensive for large datasets, but is relatively far cheaper than performing repeat runs using xarray. Slightly increased disk space requirements are incurred. We provide data for download as parquet files, but Apache Spark is not currently integrated with the framework and this data will need to be decompressed before use.
@@
We provide data chips at a fixed spatial size of $448\times 448$ m. While we provide the facility to re-tile data straightforwardly at different sizes with geetiles and sartiles, this process is computationally expensive when tiling over large spatial areas.
@@
While we provided some baseline results (Appendix D) using our framework and data, we did not provide a benchmarking framework under which models could be compared. For example, we made no assertion as to whether models should be evaluated on chips with missing channels — we chose to fill any missing data with a dummy value of 00. We also did not include European data in our test set. Although comparisons could be made by copying our approach exactly, we encourage domain experts to evaluate models according to the needs of their particular application.
Appendix G Licenses
The licenses under which each component of M3LEO was originally distributed are listed in Table G. We distribute our framework and data under the Creative Commons BY-SA 4.0 license..